OpenStudy (anonymous):

Find the equation of the tangent line to the curve f(x)=x^2+5x at x=2

OpenStudy (anonymous):

You are familiar with derivatives I hope? :-)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I am trying but I am really bad at math!

OpenStudy (anonymous):

I'm not even sure where to start with this one..

OpenStudy (anonymous):

First, you'll want to take the derivative of f(x) Second, plug in x=2 into the new function f'(x) you just found. Third, take that slope (call it 'm') and put into the point-slope formula which looks like this: \[y-y_0=m(x-x_0)\] Fourth, put x=2 into your original function f(x), to find x_0 and y_0, your coordinates of your tangential point where the tangent line just touches your parabola Finally, put in everything into your point-slop formula and make it look like y = mx + b, tada! If you get stuck give a shout out for help :D But see if you can do it first, I think you can

OpenStudy (anonymous):

The derivative gives you a function of the slope of your original function in terms of the input variable (x) in this case.

OpenStudy (anonymous):

would the derivative be 2x+5

OpenStudy (anonymous):

did you learn the shortcuts of taking derivatives yet? or are you supposed to find the derivative using the definition still?

OpenStudy (anonymous):

I don't think that I have learned the shortcuts, do you mean like the power rule and things.

OpenStudy (anonymous):

\[f'(x) = \frac{d}{dx} x^2 + 5x = 2x^{2-1}+5x^{1-1}=2x^1+5x^0=2x+5\]

OpenStudy (anonymous):

Does that look familiar?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What you'll do is input 2 for x into that function

OpenStudy (anonymous):

Read the steps I posted above :-P

OpenStudy (anonymous):

y=9?

OpenStudy (anonymous):

yes, but more specifically that's 'm', your slope of the tangent line

OpenStudy (anonymous):

That's the slope at that exact moment on your parabola at the point where x=2

OpenStudy (anonymous):

hint for step #4: \[x_0 = 2\]

OpenStudy (anonymous):

y'=9 technically, sorry minor detail

OpenStudy (anonymous):

woah, woah, woah, im now completely confused

OpenStudy (anonymous):

I dont know where we are putting everything in this equation: y−y0=m(x−x0)

OpenStudy (anonymous):

You're doing fine! You found the slope, now put that sucker into your point slope formula and go find your x_0 and y_0, your coordinates

OpenStudy (anonymous):

@jeanniebean , In your original function, what do you get as your output when you put in 2 for x?

OpenStudy (anonymous):

ok so y=mx+b is y=2(2)+5 ?

OpenStudy (anonymous):

Hold on, just answer the question I just asked for a moment. You'll be answering your own question

OpenStudy (anonymous):

14

OpenStudy (anonymous):

y=x^2+5x @ x=2 ... y=???

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

So isn't the point (2,14) our coordinate?

OpenStudy (anonymous):

oh!!!

OpenStudy (anonymous):

\[(x_0,y_0)\]

OpenStudy (anonymous):

plug & chug now into the point-slope formula. Go go go banana! :D

OpenStudy (anonymous):

y−y0=m(x−x0) = y-14=m(x-2)

OpenStudy (anonymous):

You have 'm' already remember? :P

OpenStudy (anonymous):

is that 9?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

so y-14=9(x-2)

OpenStudy (anonymous):

Distribute the 9, add 14 to both sides to move it over to the right, and you have....? (answer)

OpenStudy (anonymous):

how do you distribute it to both sides?

OpenStudy (anonymous):

oh never mind... sorry read it wrong

OpenStudy (anonymous):

I was going to say! lol

OpenStudy (anonymous):

:-P sheesh (jk jk)

OpenStudy (anonymous):

So do you have you answer now?

OpenStudy (anonymous):

y-19x - 4 ?!?

OpenStudy (anonymous):

*sigh* no... \[y-14=9(x-2)\] \[y-14=9x-18)\] \[y=?\]

OpenStudy (anonymous):

9 not 19... y=9x - 4 ?!?

OpenStudy (anonymous):

Tada!!! You did it! gj

OpenStudy (anonymous):

yay thanks soo much!

OpenStudy (anonymous):

those two intersect at EXACTLY (2,14) where the slope is 9 for both